For Problems 1 to 12, if it is appropriate to do so, use the normal approximation to the binomial to calculate the indicated probabilities

n= 300,p= 0.95,P(280 <X< 290)
Answer #1

If X ~Bin (n,p )

then , normal approximation to binomial states that , if np>10 and n(1-p) >10 thenX approximately follows a normal distribution with mean = np and variance = np(1-p) .

Here np = 285 and n(1-p) = 15 . So condition satisfies .

So, we can say approximately X~N( 285 , 3.7752)

So,

P[280<X<290] \\\\= P[\frac{280-285}{3.775}<\frac{X-285}{3.775} < \frac{290-285}{3.775}] \\\\= P[-1.3245 <Z<1.3245] \\\\= 2 \Phi(1.3245)-1 \\\\= 0.8147